### v1

# PromptHSI: Universal Hyperspectral Image Restoration Framework for Composite Degradation

Abstract
Recent developments in All-in-One (AiO) RGB image restoration and prompt learning have enabled the representation of distinct degradations through prompts, allowing degraded images to be effectively addressed by a single restoration model. However, this paradigm faces significant challenges when transferring to hyperspectral image (HSI) restoration tasks due to: 1) the domain gap between RGB and HSI features and difference on their structures, 2) information loss in visual prompts under severe composite degradations, and 3) difficulties in capturing HSI-specific degradation representations through text prompts. To address these challenges, we propose PromptHSI, the first universal AiO HSI restoration framework. By leveraging the frequency-aware feature modulation based on characteristics of HSI degradations, we decompose text prompts into intensity and bias controllers to effectively guide the restoration process while avoiding domain gaps. Our unified architecture excels at both fine-grained recovery and global information restoration tasks. Experimental results demonstrate superior performance under various degradation combinations, indicating great potential for practical remote sensing applications. The source code will be publicly released.
Corresponding author, email: cchsu@gs.ncku.edu.tw
1 Introduction
Figure 1: The All-in-One and universal HSI restoration framework with text-prompt guidance, named PromptHSI, for composite degradation Specifically, informative textual features extracted by pre-trained vision-language model are fed into adapter and then decomposed into an intensity and bias embeddings for degradation-specific restoration with effective cascade guidance.
Hyperspectral images encompass a greater number of spectral bands than RGB images, offering extensive spectral details [44]. Because materials reflect different spectral bands, each substance can have a unique signature for material identification [1, 37, 44, 32]. HSI can capture these valuable information and has broad application value and potential. However, because of the limitation of the device and bandwidth, observed HSIs are usually damaged by different types of degradation from camera sensing and transmission. Thus, HSI restoration is a crucial pre-processing step to recover degraded pixels from flawed observations.
Previous works tended to be based on HSI structures, low-rank sparsity prior [43, 11, 42] for specific degradations such as Gaussian noise [25, 49], stripe effect [24, 22], spatial blurring [33, 19], and cloud occlusion [35, 48]. However, in real-world scenario, HSIs are easily prone to multiple types of degradation, causing these methods to be limited. When HSIs are subject to multiple degradation, some methods based on deep unfolding network [4, 34] or optimization [29, 28, 52, 42] perform significantly below expectations and lack flexibility to use for real-world application.
With the development of AiO RGB image restoration methods [39, 5, 31, 10, 13], which can address multiple degradations in a single network, there is a potential to resolve this situation. However, due to the significant domain gap between the HSI and RGB images, these AiO methods cannot be transferred to HSI restoration. In particular, the visual-prompt approaches [39, 5] are influenced by severe and diverse corruptions. Furthermore, while text-prompts [10, 13] have emerged as an effective approach for controlling All-in-One restoration models, the textual feature representations encoded by widely-used vision-language models (VLMs) pre-trained on RGB-text datasets, such as CLIP [40], exhibit significant domain gaps with HSI features. Given these limitations, leveraging prompt learning frameworks for HSI processing remains an open challenge.
Considered as a whole, the essential questions are Can we leverage prompt learning and vision-language model (VLM) frameworks to address critical challenges in hyperspectral image processing for practical applications, such as achieving composite degradation restoration or controllable restoration through a single unified model? The response is affirmative.
To meet these challenges, we propose PromptHSI, a frequency-aware and prompt-guided framework tailored for HSI’s unique spectral structure and degradation types. Unlike AiO RGB image restoration methods, it ensures spectral fidelity and adaptive degradation response through novel feature modulation strategies. Our universal restoration architecture effectively handles both fine-grained recovery (e.g. spatial-spectral super-resolution) and global information restoration (e.g. decloud, inpainting) within a single model; a frequency-aware feature modulation strategy that decomposes text-prompts into intensity and bias embeddings based on frequency-domain analysis, dispelling the domain gap between VLM and HSI features. In summary, our main contributions can be listed as follows:
  • We present the first universal HSI restoration framework capable of handling composite degradations. Unlike existing HSI restoration methods that suffer from performance drops in composite degradations, our approach maintains consistent high performance across various degradation combinations.
  • We propose a novel text-to-feature modulation mechanism that effectively leverages frequency-domain characteristics of HSI degradations. By decomposing textual features into intensity and bias controller embeddings, our method achieves superior control over the restoration process, outperforming state-of-the-art AiO RGB image restoration methods on composite degradation.
  • Extensive experiments demonstrate that our method not only excels at HSI restoration under various degradation combinations but also exhibits strong generalization to unseen degradation types through flexible text-prompt control. The framework’s effectiveness is thoroughly validated through comprehensive ablation studies and comparisons with existing approaches.
Figure 2: Illustration of the degraded HSI and corresponding description synthesis pipeline. Each degradation is activated with probability $p$. Note that we set $p$ = $0.5$ for all experiments in this paper. To ensure diversity and realistic in the synthetic dataset, each activated degradation selects and applies one of its sub-degradations to the inputted HSI. (More detailed information and setting can be found in Supplementary.)
2 Related Works
Figure 3: Left, displayed from top to bottom, are HSIs with specific degradation alongside the Fourier spectra of the residual images, which are derived from subtracting the degraded images from the ground truth. Right, the diagram presents the mean values of the Fourier spectra based on the square length indicated on the x-axis. It is evident that varying degradations focus on distinct frequency subbands.
2.1 Hyperspectral Image Restoration
Conventional approaches are based on the structural properties of HSI, such as low-rank and sparse priors [43, 11]. Optimization-based methods [52, 36, 42, 6] were initially used for HSI restoration. With the progress of deep learning, HSI can be effectively recovered by imposing appropriate constraints and suitable design for task-specific restoration. Unlike RGB image restoration, HSI restoration considers not only global information aggregation but also spatial-spectral correlation at locality. Several works against different types of restoration model, such as denoising [24, 38, 15] for Gaussian noise and impulse noise, inpainting [6, 52, 29] for stripe effect, and spatial blurring [20, 26, 18] or limited spectral resolution [14, 2, 3].
However, due to multiple or composite degradations that jointly affect structural information and fine-grained details of a given HSI, restoration methods are easily compromised, consistently leading to suboptimal performance.
2.2 All-in-One RGB Image Restoration
A recently emerging paradigm in RGB image restoration involves designing single models to restore different types of degradation [12, 7, 23, 39, 10, 5, 13]. Current techniques, as a type of multi-task learning, utilize various frameworks to manage inputs and outputs for multiple tasks with prompting methods, facilitating effective information sharing between them and corresponding degradation. Specifically, AirNet [23], the first proposed AiO image restoration method, employed contrastive learning to train the encoder to learn degradation representations and restoration with a single network. PromptIR [39] introduced the visual prompt for the degradation perception and information integration within the decoding process without a priori needs. InstructIR [10] followed the text-prompt as instruction to guide the network, archiving great results with controllability. HAIR [5] is based on the hypernetwork, which aims to generate network parameters for corresponding degradations. Recently, OneRestore [13] has extended AiO restoration from multiple degradation to composite degradation, which can address the various corruption even they exists in a same image simultaneously.
Although these methods are effective, these solutions are limited in HSIs. Due to domain gap and composite degradation, it is difficult to effectively learn the compact representation of degradation and perform effective restoration.
3 Preliminary and Motivation
3.1 Unified HSI Restoration Network Design
For comprehensive restoration using a single network, employing an architecture suitable for various degradations is crucial [8, 47]. In RGB image restoration, two classic architectures predominate: ❶ U-shape encoder-decoder [46, 45] and ❷ plain residual-in-residual [27, 17, 50]. The former excels at capturing general representations and information, being widely used in denoising, deblurring, and inpainting, but pays less attention to fine details or repeat patterns. The latter specializes in capturing detailed spatial features at the fine-grained level, serving as the backbone for super-resolution tasks [9, 41, 51].
Previous methods have overlooked the compatibility between these two restoration models, and current AiO restoration approaches [39, 5, 13] typically focus on the former rather than integrate the joint strength of both types of network design, limiting their versatility. Besides, this limitation also extends to HSI restoration models [33, 19, 21]. Therefore, we integrate a plain residual-in-residual block into the U-net cascade decoder in the PromptHSI to maintain the advantages of both designs without compromising. The versatility of the network can be ensured accordingly.
3.2 Degradations Analysis in Frequency Domain
Unlike RGB image restoration, HSI restoration requires careful consideration of spatial-spectral correlations across hundreds of bands. Our key insight comes from analyzing how different types of degradation manifest in the frequency-domain through Fourier analysis. Given a clean HSI $I_{\text{clean}}$, its Fourier transform can be expressed as:
$F(I_{\text{clean}})=F_{s}(I_{\text{clean}})+iF_{c}(I_{\text{clean}})$ (1)
where $F_{s}$ and $F_{c}$ represent the sine and cosine components respectively. Through empirical analysis, as shown in Figure 3, we observed that different types of degradation exhibit distinct patterns in how they affect these frequency components.
Specifically, the relationship between degraded HSI $I_{\text{degraded}}$ and clean HSI $I_{\text{clean}}$ in frequency-domain can be characterized by an affine transformation:
$F(I_{\text{degraded}})=\hat{\lambda}\odot F(I_{\text{clean}})+\hat{\mu}=(1+\lambda)\odot F(I_{\text{clean}})+\mu$ (2)
where $\odot$ denotes element-wise multiplication, $\lambda$ and $\mu$ are degradation-specific parameters that: $\lambda$ represents the scaling factor that modulates frequency amplitudes, $\mu$ is the bias term that shifts the frequency distribution.
Different types of degradation show characteristic patterns in $\lambda$ and $\mu$, such as ❶ Blur mainly affects high frequencies: $\lambda_{high}$ $\ll$ $\mu_{low}$. ❷ Noise adds uniform perturbations: $\mu_{noise}$approx constant $c$. ❸ Missing information creates structured patterns in both $\lambda$ and $\mu$.
4 Proposed Method
Figure 4: The architecture of the proposed PromptHSI, which is designed to tackle the composite degradations within a single universal model. First, we design a network based on U-net-like encoder-decoder architecture [46, 45] for capturing global spatial-spectral information, while combining Residual-Dense Group (RDG) and Fast-Residual-Dense Block (FRDB) presented in DRCT [17] and RTCS [16], respectively, in the proposed PGFAB for spatial-spectral enhancement. Afterwards, text-prompt-guidance and frequency-aware feature map modulation is proposed to jointly better capture the representation of composite degradation with controllability.
Overview. Figure 4 illustrates the overall architecture of PromptHSI, which consists of three key components designed to address the challenges in universal HSI restoration: ❶ A unified U-shape network which contains Spatial-Spectral Aware Feature Extraction Block (SSAEFB) and Propmpt-guided Feature Aggregation Block (PGFAB) that effectively captures both enhanced spatial-spectral fine-grained details and global contextual information from degraded HSI; ❷ A Prompt-Guided Feature Modulation (PGFM) that effectively bridging the domain gap between textual features and HSI features through frequency-domain modulation, while avoiding direct cross-domain feature alignment and information loss.
4.1 Hyperspectral Image Restoration Pipeline
Our restoration pipeline consists of two key components: a SSAEFB for efficient feature extraction, and a PGFAB for controllable restoration with text-prompt guidance.
The encoder exploits the low-rank and sparse prior of HSIs through depth-wise separable convolution, effectively reducing computational complexity while preserving spectral characteristics. PGFAB then utilizes dual branches for spatial and spectral enhancement, with cross-attention mechanisms facilitating comprehensive feature integration.
Spatial-spectral-aware encoder. In conventional RGB restoration and other vision tasks, standard $3\times 3$ convolutions are widely considered to be sufficiently versatile and effective feature encoders. However, considering the low-rank characteristics of HSI, where substantial redundancy exists between different spectral bands, utilizing standard convolutions may reference unnecessary information from distant spectral bands, leading to suboptimal performance. Therefore, encoders based on lightweight convolutions prove sufficiently effective while further reducing parameter size to meet practical application requirements.
Specifically, for the input HSI $I_{\text{degraded}}$, we first interpolate the spectrally discontinuous regions (caused by the stripe effect) to facilitate the learning process. Subsequently, we apply standard convolution layers to obtain shallow feature maps.
Afterwards, SSAFEBs is based on a depth-wise separable convolution to extract features while inherently taking advantage of this low-rank prior, thereby reducing computational complexity without compromising performance [16].
Propmpt-guided feature aggregation block. The features captured by SSAFEB are combined with the informative text-prompt guidance from PGFM (which will be discussed in the next subsection). Using effective guidance, the network is allowed to dynamically activate the corresponding weight for HSI restoration according to different types of degradation. To capture detailed information, PGFAB utilizes two distinct branches: one focuses on spatial dimension and the other on spectra. These branches are eventually integrated for feature aggregation.
Spatial and spectral feature enhancement. The output of PGFM is first improved by consecutive RDG [17] and FRDB [16] to improve fine-grained details. Then these features are interacted by cross-attention layer, which aims to comprehensively integrate the spatial-spectral feature information from different branches across all dimensions. Specifically, it can be expressed as follows:
$\{Q_{\alpha},K_{\alpha},V_{\alpha}\}=\mathcal{F}^{\alpha}_{qkv}(F_{\alpha}),\quad\{Q_{\beta},K_{\beta},V_{\beta}\}=\mathcal{F}^{\beta}_{qkv}(F_{\beta}),$ (3)
where $F_{\alpha},F_{\beta}$ denote features from the spatial and spectral enhancement block. Subsequently, we exchange the queries $Q$ of two branches for spatial-spectral interaction:
$F_{f}^{\beta}=softmax(\frac{Q_{\alpha}K_{\beta}}{d_{k}})V_{\beta},\quad F_{f}^{\alpha}=softmax(\frac{Q_{\beta}K_{\alpha}}{d_{k}})V_{\alpha},$ (4)
where $d_{k}$ is the scaling factor. Subsequently, we concatenate the results obtained by the cross-attention through $F_{f}^{\gamma}=Concat({F_{f}^{\beta},F_{f}^{\alpha}})$ to obtain the integrated features.
$\{Q_{\gamma},K_{\gamma},V_{\gamma}\}=\mathcal{F}^{\gamma}_{qkv}(F_{f}^{\gamma}),\quad\hat{F_{f}^{\gamma}}=softmax(\frac{Q_{\gamma}K_{\gamma}}{d_{k}})V_{\gamma},$ (5)
Finally, we employ spectral-spatial self-attention for feature co-enhancement to improve representation capability.
Through a dual-branch architecture, we integrate text-prompt guidance effectively across both spatial and spectral dimensions. By incorporating plain-residual-in-residual [27, 17, 16] modules within the U-shape cascade decoder [46, 8], we enhance fine-grained image restoration while integrating global information, thereby effectively addressing various degradations for HSI restoration.
Table 1: Overall performance evaluation for composite degradation and complexity comparison. To simulate composite degradations, we maintain a fixed $p$ of 0.5 for triggering individual degradation types, as demonstrated in Figure 2, for all experiments. Note that Cr. means Controllable or Not. As for the complexity parts, M and G indicate $10^{6}$ and $10^{9}$, respectively. The best, second-best and third-best results are highlighted in bold-red, bold-blue, and bold-green, respectively.
[…]
Input - - - - $\times$ 15.2454 17.7463 0.0643 33.7220 - -
HSIDwRD [49] ICCV 2021 Denoising HSI $\times$ 20.6270 12.7891 0.0364 97.3601 23.62M 5054.16G
SERT [25] CVPR 2023 Denoising HSI $\times$ 16.1657 15.9853 0.0348 29.2053 2.74M 176.06 G
SRDNet [33] TGRS 2024 Super-Res. HSI $\times$ 19.0520 11.9590 0.0320 16.5861 1.86M 3771.80G
SNLSR [19] TIP 2024 Super-Res. HSI $\times$ 17.8785 20.4389 0.0445 19.1267 1.75M 14.20G
One-to-One SGNet [35] ISPRS 2022 Decloud HSI $\times$ 18.0694 12.9476 0.0359 17.9773 0.54M 57.39G
AACNet [48] TGRS 2023 Decloud HSI $\times$ 17.6279 14.5962 0.0324 26.2115 2.76M 171.23G
SST [24] AAAI 2023 Inpainting HSI $\times$ 18.9792 10.9860 0.0244 26.9934 4.30M 281.72G
HSDT [22] ICCV 2023 Inpainting HSI $\times$ 18.0665 11.3451 0.0372 68.4451 0.51M 67.55G
AirNet [23] CVPR 2022 - RGB $\times$ 23.6647 7.5970 0.0233 16.4194 6.07M 245.86G
All-in-One PromptIR [39] NeurIPS 2023 - RGB $\times$ 17.2667 9.6437 0.0405 29.2768 33.18M 132.07G
for specific HAIR [5] arXiv 2024 - RGB $\times$ 20.5779 14.4329 0.0275 24.4211 7.86M 42.51G
InstructIR [10] ECCV 2024 - RGB $\checkmark$ 20.3106 14.5727 0.0306 27.5454 16.02M 17.28G
All-in-One OneRestore [13] ECCV 2024 - RGB $\checkmark$ 17.3321 15.0320 0.0271 52.1162 5.85M 12.05G
for composite PromptHSI (Ours) - - HSI $\checkmark$ 26.4432 6.0601 0.0185 8.3566 26.14M 135.83G
Figure 5: Visualization results for HSI restoration methods. top: the reconstructed HSI, bottom: the residual image derived from Ground-truth and reconstructed result. Most of state-of-the-arts HSI restoration methods fail to deal with composite degradations.
4.2 Prompt-Guided Feature Modulation
Directly using text features from VLMs for HSI restoration faces several critical challenges. Firstly, VLM features are trained on RGB-text pairs, lacking understanding of spectral characteristics and capturing complex spectral degradations. Secondly, their statistical distribution are mismatched because VLM features follow RGB image statistics, but HSI has essentially different data distributions, and direct adaptation requires complex domain translation.
Instead of directly bridging the HSI-text domain gap, our proposed PGFM addresses the domain gap by decomposing the problem into two easier sub-tasks. ❶ frequency-pattern matching with task descriptor and ❷ distribution adjustment by novel frequency-aware feature modulation mechanism. This strategy aim to selectively enhance spectral fidelity across frequency bands, thereby overcoming the limitations of text-guidance for avoiding direct cross-domain alignment between text features extracted from VLMs pre-trained on RGB-text domains and HSI.
Task descriptor generation. Given a text prompt $T_{text}$ that provides the corresponding degradation description to guide the HSI restoration network to obtain the specified restoration result. In this study, we employ CLIP [40], which integrates vision and language information and performs effectively in text feature extraction. We freeze the pre-trained CLIP to maintain great linguistic consistency and add additional MLP layer $\Phi_{t}^{adpt.}$ as the feature adapter for pattern matching. With {}$\{\cdot\}_{e}$ denoting the frozen weights, this process can be expressed as:
$\footnotesize F_{text}=\Phi_{t}^{adpt.}(\{\mathcal{F}^{CLIP}_{text}\}_{e}(T_{text})),$ (6)
where $F_{text}$ denotes the task descriptors. With the task descriptors available, the next challenge is how to effectively incorporate this guidance into the HSI restoration process.
Frequency-aware feature modulation. Based on the frequency domain analysis, as illustrated in Section 3.2, we design our feature modulation mechanism to mimic how degradations affect different frequency components. Given an input feature map $F_{f}$, we first decompose it into frequency components:
$F_{f}=F_{f}^{l}+F_{f}^{h},$ (7)
where $F_{f}^{l}$ and $F_{f}^{h}$ represent low and high frequency components, obtained through frequency-domain filtering.
Afterward, we design the MLP layer $\Psi_{m}^{i}$ to extract this connection with the intensity and bias embeddings for feature modulation. Therefore, it can be obtained:
$\lambda_{m}=\Psi_{m}^{I}(F_{text}),\ \mu_{m}=\Psi_{m}^{II}(F_{text}),$ (8)
where $\Psi_{m}^{I}$ and $\Psi_{m}^{II}$ are the chunk operations of $\Psi_{m}$ to form the intensity and bias controller embeddings. To address the spectral consistency challenge in HSIs, our frequency-aware feature modulation adopts an affine transformation model tailored for spectral data, formulated as:
$\hat{F}_{f}=(1+\lambda_{m}^{l})\odot F_{f}^{l}+(1+\lambda_{m}^{h})\odot F_{f}^{h}+\mu,$ (9)
where Intensity controller $\lambda$in [$\lambda_{m}^{l}$, $\lambda_{m}^{h}$] modulates the amplitude of high-frequency components to address noise, while Bias controller $\mu$ shifts the low-frequency baseline to compensate for blurring and band-missing effects. This design ensures frequency-specific adjustments that RGB AiO frameworks lack.
To validate the efficacy of this mechanism, we performed ablation studies demonstrating improved spectral fidelity and PSNR consistency across spectral bands, underscoring the need for spectral domain modulation to manage complex HSI degradations effectively.
4.3 Loss Function
Given the wide range of spectra inherent in HSI, we utilize only the $\ell_{1}$ loss for training. However, this strategy may not adequately capture the variability across different bands. Consequently, the $\ell_{1}$ loss might inadvertently focus on minimizing bands with larger values, affecting the spectral integrity of specific bands.
Spectral angle mapper loss. SAM metric is also critical for material identification and less sensitive to intensity variations. The $\ell_{\text{SAM}}$ loss prioritizes angular discrepancies over magnitudes, offering a reliable metric to evaluate the spectral fidelity of the reconstructed HSI.
$\ell_{\text{SAM}}=\frac{1}{N}\sum_{n=1}^{N}\cos^{-1}\left(\frac{I_{n}^{T}I^{*}_{n}+\epsilon}{\|I_{n}\|_{2}\cdot\|I_{n}^{*}\|_{2}+\epsilon}\right),$ (10)
where $I_{n}$ and $I^{*}_{n}$ represent the $n$-th spectral vector of the ground-truth and the reconstructed HSI, respectively. The term $\epsilon$ is included to avoid divergence.
Stationary wavelet transform loss. Wavelet sub-band alignment is able to manage various aspects of information encoded in HSIs. On the other words, it facilitates more detailed supervision in the frequency-domain.
$\ell_{\text{SWT}}=\sum_{j}\lambda_{j}\big{\|}\text{SWT}(I)_{j}-\text{SWT}(I^{*})_{j}\big{\|}_{1},$ (11)
where $\lambda_{j}$ represents the count of sub-bands decomposed via stationary wavelet transformation.
Band-wise MSE loss. The band-wise disparity helps stabilize the network optimization process and is advantageous to achieve convergence, balancing restoration quality across all spectra. The Band-wise MSE loss can be simply defined as follows:
$\ell_{\text{BMSE}}=\frac{1}{N}\sum_{n=1}^{N}{\|I_{n}-I_{n}^{*}\|_{2}},$ (12)
Total loss. Ultimately, the total loss function for training the proposed PromptHSI $\ell_{\text{total}}$ can be expressed as:
$\ell_{\text{total}}=\lambda_{1}\ell_{\text{1}}+\lambda_{2}\ell_{\text{SAM}}+\lambda_{3}\ell_{\text{SWT}}+\lambda_{4}\ell_{\text{BMSE}},$ (13)
where $\lambda_{1,2,3,4}$ serve as the weighting coefficients to equilibrate $\ell_{1}$, $\ell_{\text{SAM}}$, $\ell_{\text{SWT}}$, and $\ell_{\text{BMSE}}$, set at 1, 0.1, 0.01, 0.01 throughout all experiments, respectively.
5 Experimental Results
Figure 6: Comparison for AiO RGB image restoration methods. For the text-prompt-based methods [10, 13] and our PromptHSI, we employ the standard text-prompts presented in this study for training network and inference for related experiments.
Figure 7: Comparison of t-SNE visualization results for the degradation embeddings of the proposed PromptHSI trained on single degradation. (a) for multiple degradation, (b) for composite degradatation. The proposed PromptHSI have demonstrated its discriminative capability in identifying different degradation even unseen degradation and its combination during testing-time.
Table 2: Ablation of integration strategy.
Strategy PSNR / SAM / RMSE / EGRAS
Intensity-only 26.0070 / 6.3636 / 0.0187 / 9.7608
Bias-only 25.6384 / 6.3963 / 0.0190 / 12.4992
Weighted-sum 25.7284 / 6.3561 / 0.0188 / 14.3605
Intensity & Bias 26.4432 / 6.0601 / 0.0185 / 8.3566
Table 3: Effect of prompting method.
Prompt type Cr. PSNR / SAM / RMSE / EGRAS
RGB to VLM $\times$ 24.1273 / 7.6236 / 0.0257 / 14.5208
Hybrid to VLM $\checkmark$ 25.2385 / 6.8974 / 0.0191 / 11.4359
PGB & PIM [39] $\times$ 23.0567 / 9.4862 / 0.0287 / 24.5575
Text to VLM $\checkmark$ 26.4432 / 6.0601 / 0.0185 / 8.3566
Table 4: Ablation of network component.
Without PSNR / SAM / RMSE / EGRAS
Full model 26.4432 / 6.0601 / 0.0185 / 8.3566
PGFM 25.4905 / 6.6348 / 0.0205 / 13.2072
adapter 25.8686 / 6.3561 / 0.0188 / 14.3605
enhancement 23.7511 / 7.9918 / 0.0262 / 16.0007
Experiment Setup. This study employed the AVIRIS dataset provided by Lin et al. [30, 18], an extensive HSI dataset, with numerous bands spanning diverse geographical areas for experiments. It includes 2,078 HSIs, which were randomly divided into training, validation, and testing sets. The training set consisted of 1,678 images and both the validation and testing sets contained 200 images each. As illustrated in Supplementary, all clean HSI was pre-processed to synthesize the composite degraded HSIs and corresponding text descriptions.
5.1 Quantitative Results and Analysis
To evaluate the performance of PromptHSI, we compared it with eight other supervised HSI restoration methods for specific degradation, including HSIDwRD [49], SERT [25], SRDNet [33], SNLSR [19], SGNet [35], AACNet [48], SST [24], and HSDT [22] in the synthesized AVIRIS dataset with composite degradations. Because there is no AiO HSI restoration method, we compare four AiO RGB image restoration methods in the experiments: AirNet [23], PromptIR [39], HAIR [5], InstructIR [10], and OneRestore [13]. The model performance was measured using the four metrics, including PSNR, SAM, RMSE, and ERGAS. Table 1 presents the quantitative results.
Ours vs. Task-specific HSI restoration methods. All peer-methods fail to effectively recover the degraded HSIs due to their ability to learn different or unseen degradations. In contrast, PromptHSI can restore better even without text-prompt, as shown in Table 4, which exhibits the capability of our unified restoration network design.
Ours vs. All-in-One RGB restoration methods. Unlike AiO RGB image restoration methods that rely on uniform feature modulation, our PromptHSI framework’s frequency-aware modulation dynamically tailors restoration to each spectral band, thereby achieving cross-band consistency crucial for HSI applications.
In general, our PromptHSI exceeds other AiO methods. Firstly, most of existing methods struggle to handle composite degradations. As discussed previously, both data and features extracted by pre-trained CLIP [40] exhibit domain gaps with HSI. Therefore, these models are not effective in transferring to AiO HSI restoration.
Specifically, OneRestore [13] incorporated task description features into its encoder block, which hindered its capacity to accurately capture the spatial-spectral correlation in HSI because of the domain gap with HSI features, resulting in decoding failures. Moreover, it lacks considerations for tasks requiring high-frequency detail reconstruction, potentially limiting its effectiveness in the spatial or spectral super-resolution task.
5.2 Ablation Studies and Analysis
Figure 8: Visualization results for the controllability. Top, for single-specific degradation, the plots in the first row represent the degraded HSI, while the second row shows residual maps, obtained from the reconstructed HSI, with or without the guidance provided by the text-prompt. Bottom, in the case of composite degradation, the top row of plots depicts the recovered HSI, while the residual maps are obtained from the reconstructed HSI using the specified text-prompt guidance with the ground-truth.
The aforementioned experimental results demonstrate the effectiveness of our proposed PromptHSI framework under composite degradations. We further conduct ablation studies to investigate the impact of different components and strategies.
Integration strategy of feature-map modulation. As shown in Table 4, our proposed frequency-aware feature modulation demonstrates superior effectiveness compared to the weighted-sum manner, where we implemented an MLP layer for dynamic weighted averaging of inputs and outputs. The intensity-only configuration outperforms the bias-only, with a notably larger advantage in the ERGAS metric, suggesting that this integration strategy effectively guides network restoration across most scenarios.
Visual-prompt or text-prompt. The type of prompting is crucial, as it determines its ability to recognize various degradations and effectively guide and control model behavior. In the first configuration, we select RGB bands from HSIs as visual prompts, encoding them through CLIP [40] and using the proposed PGFM to guide the model. Secondly, we encode both RGB bands and textual prompts to generate hybrid task descriptors for guidance. In the final configuration, we remove PGFM and use the PIM and PGM block proposed in PromptIR [39] for the generation of visual prompts and interaction of features.
The experimental results are presented in Table 4. Due to multiple or composite degradations in the spatial-domain leading to serious information loss, it became difficult to obtain compact task descriptors or valid visual prompts, ultimately resulting in diminished prompt effectiveness and compromised restoration. The use of textual features helps to circumvent these issues and achieves robust prompting guidance through semantic information.
Ablation of network component. To understand the impact of the network components, we performed ablation studies by removing PGFM, adapter and spatial spectral enhancement from the network, with results shown in Table 4. Even without textual features, our proposed PromptHSI outperforms other peer methods in handling composite degradation, demonstrating the effectiveness and universality of our network architecture. The adapter proves crucial in bridging the domain gap between the VLM and HSI features via pattern matching, leading to improved global reconstruction performance, particularly evident in the ERGAS metric. Lastly, we can observe that there is a drop with or without spatial-spectral enhancement, which represents that feature enhancement is essential for details recovery within cascade decoder.
These results demonstrate PromptHSI’s superior capability in maintaining spectral fidelity across multiple degradation types, highlighting the necessity of frequency-aware modulation, particularly when compared to AiO RGB methods that exhibit substantial performance degradation in the HSI domain. In summary, PromptHSI offers a universal and controllable solution for addressing composite degradation in real-world applications.
6 Conclusion
In this paper, we present PromptHSI, the first universal framework for hyperspectral image restoration that effectively addresses composite degradations, which conventional HSI restoration methods have failed to handle. Unlike existing All-in-One RGB image restoration approaches that struggle with domain gaps between RGB and HSI features and structures, our method is specifically designed to address HSI-specific attributes and commonly encountered degradation patterns. Then, our method is controllable through text-prompt guidance, based on a novel frequency-aware modulation strategy that effectively bridges the gap between text and HSI through frequency pattern matching and distribution adjustment in the frequency-domain. Our approach demonstrates superior performance across various types of degradation and their combinations, significantly outperforming existing methods.

### v3

@@
-PromptHSI: Universal Hyperspectral Image Restoration Framework for Composite Degradation
+PromptHSI: Universal Hyperspectral Image Restoration with Vision-Language Modulated Frequency Adaptation
 Abstract
-Recent developments in All-in-One (AiO) RGB image restoration and prompt learning have enabled the representation of distinct degradations through prompts, allowing degraded images to be effectively addressed by a single restoration model.
-However, this paradigm faces significant challenges when transferring to hyperspectral image (HSI) restoration tasks due to: 1) the domain gap between RGB and HSI features and difference on their structures, 2) information loss in visual prompts under severe composite degradations, and 3) difficulties in capturing HSI-specific degradation representations through text prompts.
-To address these challenges, we propose PromptHSI, the first universal AiO HSI restoration framework.
-By leveraging the frequency-aware feature modulation based on characteristics of HSI degradations, we decompose text prompts into intensity and bias controllers to effectively guide the restoration process while avoiding domain gaps.
-Our unified architecture excels at both fine-grained recovery and global information restoration tasks.
-Experimental results demonstrate superior performance under various degradation combinations, indicating great potential for practical remote sensing applications.
-The source code will be publicly released.
-Corresponding author, email: cchsu@gs.ncku.edu.tw
+Recent advances in All-in-One (AiO) RGB image restoration have demonstrated the effectiveness of prompt learning in handling multiple degradations within a single model.
+However, extending these approaches to hyperspectral image (HSI) restoration is challenging due to the domain gap between RGB and HSI features, information loss in visual prompts under severe composite degradations, and difficulties in capturing HSI-specific degradation patterns via text prompts.
+In this paper, we propose PromptHSI, the first universal AiO HSI restoration framework that addresses these challenges.
+By incorporating frequency-aware feature modulation—which utilizes frequency analysis to narrow down the restoration search space—and employing vision-language model (VLM)-guided prompt learning, our approach decomposes text prompts into intensity and bias controllers that effectively guide the restoration process while mitigating domain discrepancies.
+Extensive experiments demonstrate that our unified architecture excels at both fine-grained recovery and global information restoration across diverse degradation scenarios, highlighting its significant potential for practical remote sensing applications.
+The source code is available at https://github.com/chingheng0808/PromptHSI.
+Corresponding author, email: cchsu@gs.ncku.edu.twProject page: https://chingheng0808.github.io/prompthsiP/static.html
 1 Introduction
-Figure 1: The All-in-One and universal HSI restoration framework with text-prompt guidance, named PromptHSI, for composite degradation Specifically, informative textual features extracted by pre-trained vision-language model are fed into adapter and then decomposed into an intensity and bias embeddings for degradation-specific restoration with effective cascade guidance.
+Figure 1: Performance comparison between All-in-One and HSI restoration methods on the AVIRIS dataset.
+The proposed PromptHSI offers leading performance and controllability while retaining a compact model size.
 Hyperspectral images encompass a greater number of spectral bands than RGB images, offering extensive spectral details [52].
 Because materials reflect different spectral bands, each substance can have a unique signature for material identification [1, 45, 52, 38].
-HSI can capture these valuable information and has broad application value and potential.
+HSI can capture this valuable information and has broad application value and potential.
 However, because of the limitation of the device and bandwidth, observed HSIs are usually damaged by different types of degradation from camera sensing and transmission.
 Thus, HSI restoration is a crucial pre-processing step to recover degraded pixels from flawed observations.
-Previous works tended to be based on HSI structures, low-rank sparsity prior [43, 11, 42] for specific degradations such as Gaussian noise [25, 49], stripe effect [24, 22], spatial blurring [33, 19], and cloud occlusion [35, 48].
-However, in real-world scenario, HSIs are easily prone to multiple types of degradation, causing these methods to be limited.
-When HSIs are subject to multiple degradation, some methods based on deep unfolding network [4, 34] or optimization [29, 28, 52, 42] perform significantly below expectations and lack flexibility to use for real-world application.
-With the development of AiO RGB image restoration methods [39, 5, 31, 10, 13], which can address multiple degradations in a single network, there is a potential to resolve this situation.
-However, due to the significant domain gap between the HSI and RGB images, these AiO methods cannot be transferred to HSI restoration.
-In particular, the visual-prompt approaches [39, 5] are influenced by severe and diverse corruptions.
-Furthermore, while text-prompts [10, 13] have emerged as an effective approach for controlling All-in-One restoration models, the textual feature representations encoded by widely-used vision-language models (VLMs) pre-trained on RGB-text datasets, such as CLIP [40], exhibit significant domain gaps with HSI features.
-Given these limitations, leveraging prompt learning frameworks for HSI processing remains an open challenge.
-Considered as a whole, the essential questions are Can we leverage prompt learning and vision-language model (VLM) frameworks to address critical challenges in hyperspectral image processing for practical applications, such as achieving composite degradation restoration or controllable restoration through a single unified model?
-The response is affirmative.
-To meet these challenges, we propose PromptHSI, a frequency-aware and prompt-guided framework tailored for HSI’s unique spectral structure and degradation types.
-Unlike AiO RGB image restoration methods, it ensures spectral fidelity and adaptive degradation response through novel feature modulation strategies.
-Our universal restoration architecture effectively handles both fine-grained recovery (e.g.
-spatial-spectral super-resolution) and global information restoration (e.g.
-decloud, inpainting) within a single model; a frequency-aware feature modulation strategy that decomposes text-prompts into intensity and bias embeddings based on frequency-domain analysis, dispelling the domain gap between VLM and HSI features.
-In summary, our main contributions can be listed as follows:
-  • We present the first universal HSI restoration framework capable of handling composite degradations.
-Unlike existing HSI restoration methods that suffer from performance drops in composite degradations, our approach maintains consistent high performance across various degradation combinations.
-  • We propose a novel text-to-feature modulation mechanism that effectively leverages frequency-domain characteristics of HSI degradations.
-By decomposing textual features into intensity and bias controller embeddings, our method achieves superior control over the restoration process, outperforming state-of-the-art AiO RGB image restoration methods on composite degradation.
-  • Extensive experiments demonstrate that our method not only excels at HSI restoration under various degradation combinations but also exhibits strong generalization to unseen degradation types through flexible text-prompt control.
-The framework’s effectiveness is thoroughly validated through comprehensive ablation studies and comparisons with existing approaches.
-Figure 2: Illustration of the degraded HSI and corresponding description synthesis pipeline.
-Each degradation is activated with probability $p$.
-Note that we set $p$ = $0.5$ for all experiments in this paper.
-To ensure diversity and realistic in the synthetic dataset, each activated degradation selects and applies one of its sub-degradations to the inputted HSI.
-(More detailed information and setting can be found in Supplementary.)
+To address these issues, low-rank tensor models [49, 42] are widely used in hyperspectral image restoration, relying on the assumption that the clean image $\mathbf{X}$ exhibits low-rank properties [50, 13] across spatial and spectral dimensions.
+The restoration process is often formulated as $\mathbf{X}\approx\mathbf{L}+\mathbf{S}+\mathbf{N}$, where $\mathbf{L}$ represents the underlying low-rank structure of the clean image, $\mathbf{S}$ denotes structured sparse noise (e.g., stripe noise or dead pixels), $\mathbf{N}$ represents random noise, such as Gaussian or Poisson noise.
+However, under multiple degradations, this assumption often fails.
+Structured noise components $\mathbf{S}$ introduce high-rank artifacts, increasing the observed rank: $\operatorname{rank}(\mathbf{X})\gg\operatorname{rank}(\mathbf{L})$, which makes low-rank approximations less accurate.
+Furthermore, mixed noise complicates the separation between the clean component $\mathbf{L}$ and noise components $\mathbf{S}$ and $\mathbf{N}$, often leading to over-smoothing:
+$\min_{\mathbf{L}}\|\mathcal{P}_{\Omega}(\mathbf{X}-\mathbf{L})\|_{F}^{2}+\lambda\operatorname{TV}(\mathbf{L}),$ (1)
+where $P_{\Omega}$ is a projection operator on observed pixels, and $\operatorname{TV}(\mathbf{L})$ is a total variation regularization term.
+While this enforces spatial smoothness, it often suppresses fine details, degrading spectral fidelity.
+Therefore, when multiple degradations are present, the core low-rank assumption is violated, making traditional tensor-based approaches ineffective.
+More adaptive methods that incorporate hybrid priors [34, 33, 63] or deep learning-based corrections are needed to achieve robust restoration.
+The emergence of AiO RGB image restoration methods [47, 4, 36, 10, 15, 11] has demonstrated remarkable success in handling diverse degradations within a unified network.
+These methods leverage prompt learning, either through visual prompts [47, 4, 11] or text-based prompts [10, 15], to adaptively restore corrupted images.
+However, despite their effectiveness in RGB restoration, these AiO approaches face fundamental limitations when extended to HSI restoration.
+The primary challenge arises from the inherent complexity and variability of degradation patterns in HSI data, making it difficult to generalize the same prompt-based strategies.
+To address these challenges, we propose PromptHSI, a frequency-aware, prompt-guided HSI restoration framework that overcomes the limitations of existing AiO RGB-based restoration methods.
+Unlike conventional approaches that struggle with the complexity of HSI degradation patterns, PromptHSI integrates frequency-domain priors with VLM-driven prompt learning, enabling robust and adaptive restoration across diverse degradation scenarios.
+In summary, our key contributions are summarized as follows:
+  • A Frequency-Aware Prior for Efficient Restoration.
+We introduce a structured frequency-domain analysis that models the impact of different degradations on specific frequency components.
+By leveraging Fourier-based priors, PromptHSI significantly reduces the restoration search space, making it easier for deep networks to learn effective recovery mappings.
+  • Vision-Language Modulated Prompt Learning for Controllable Restoration.
+Unlike traditional prompt-based approaches that rely on static feature modulation, PromptHSI employs VLM-guided adaptive modulation, where textual embeddings dynamically control the restoration process.
+This allows for more flexible and interpretable restoration outcomes, surpassing conventional AiO RGB methods in handling complex degradations.
+  • The First Composite Degradation Dataset for Benchmarking HSI Restoration.
+To facilitate research in HSI restoration, we construct the first standardized composite degradation dataset, incorporating cloud occlusion, blurring, noise, and spectral band loss.
+We believe this dataset will be publicly released to support future developments in hyperspectral image restoration.
+Through extensive experiments, we demonstrate that PromptHSI consistently outperforms state-of-the-art AiO RGB methods in HSI restoration tasks while offering controllable restoration via text prompts.
+Figure 2: Illustration of the degraded HSI and description synthesis pipeline.
+Each degradation is activated with probability $p$ (set to 0.5 in all experiments).
+To ensure diversity and realism, each activated degradation randomly applies one of its sub-degradations to the input HSI.
+More details are provided in the Supplementary.
 2 Related Works
-Figure 3: Left, displayed from top to bottom, are HSIs with specific degradation alongside the Fourier spectra of the residual images, which are derived from subtracting the degraded images from the ground truth.
-Right, the diagram presents the mean values of the Fourier spectra based on the square length indicated on the x-axis.
-It is evident that varying degradations focus on distinct frequency subbands.
 2.1 Hyperspectral Image Restoration
@@
 Several works against different types of restoration model, such as denoising [28, 46, 17] for Gaussian noise and impulse noise, inpainting [5, 63, 34] for stripe effect, and spatial blurring [23, 30, 20] or limited spectral resolution [16, 2, 3].
-However, due to multiple or composite degradations that jointly affect structural information and fine-grained details of a given HSI, restoration methods are easily compromised, consistently leading to suboptimal performance.
+Despite these advancements, multiple or composite degradations simultaneously affect both structural integrity and fine-grained details in HSIs, making restoration methods highly prone to performance degradation.
+Moreover, erroneous compensation effects [37, 27] further exacerbate this issue, as models may misinterpret different degradation types, leading to incorrect adjustments.
+For instance, in HSIs degraded by both noise and cloud occlusion, conventional denoising methods may mistakenly treat occluded regions as random noise, resulting in spectral distortions and suboptimal restoration outcomes.
 2.2 All-in-One RGB Image Restoration
-A recently emerging paradigm in RGB image restoration involves designing single models to restore different types of degradation [12, 7, 23, 39, 10, 5, 13].
-Current techniques, as a type of multi-task learning, utilize various frameworks to manage inputs and outputs for multiple tasks with prompting methods, facilitating effective information sharing between them and corresponding degradation.
-Specifically, AirNet [23], the first proposed AiO image restoration method, employed contrastive learning to train the encoder to learn degradation representations and restoration with a single network.
-PromptIR [39] introduced the visual prompt for the degradation perception and information integration within the decoding process without a priori needs.
-InstructIR [10] followed the text-prompt as instruction to guide the network, archiving great results with controllability.
-HAIR [5] is based on the hypernetwork, which aims to generate network parameters for corresponding degradations.
-Recently, OneRestore [13] has extended AiO restoration from multiple degradation to composite degradation, which can address the various corruption even they exists in a same image simultaneously.
-Although these methods are effective, these solutions are limited in HSIs.
-Due to domain gap and composite degradation, it is difficult to effectively learn the compact representation of degradation and perform effective restoration.
+A recently emerging paradigm in RGB image restoration designs single models to handle multiple degradations [14, 6, 27, 47, 10, 4, 15].
+These multi-task methods leverage prompting and diverse frameworks to share information between tasks and degradations.
+For example, AirNet [27] used contrastive learning to jointly learn degradation representations and restoration; PromptIR [47] introduced visual prompts for degradation perception during decoding; InstructIR [10] employed text prompts to guide the network; HAIR [4] utilized a hypernetwork to generate degradation-specific parameters; OneRestore [15] extended AiO restoration to composite degradations, addressing multiple corruptions in a single image; and AdaIR [11] transforms feature maps into the frequency domain via Fourier transform and uses an adaptive mask, guided by the degraded image, to decouple low- and high-frequency features.
+However, despite their effectiveness in RGB, these solutions are limited in HSIs.
+Due to domain gaps and composite degradation, learning a compact degradation representation and achieving effective restoration in HSIs remains challenging.
 3 Preliminary and Motivation
-3.1 Unified HSI Restoration Network Design
-For comprehensive restoration using a single network, employing an architecture suitable for various degradations is crucial [8, 47].
-In RGB image restoration, two classic architectures predominate: ❶ U-shape encoder-decoder [46, 45] and ❷ plain residual-in-residual [27, 17, 50].
-The former excels at capturing general representations and information, being widely used in denoising, deblurring, and inpainting, but pays less attention to fine details or repeat patterns.
-The latter specializes in capturing detailed spatial features at the fine-grained level, serving as the backbone for super-resolution tasks [9, 41, 51].
-Previous methods have overlooked the compatibility between these two restoration models, and current AiO restoration approaches [39, 5, 13] typically focus on the former rather than integrate the joint strength of both types of network design, limiting their versatility.
-Besides, this limitation also extends to HSI restoration models [33, 19, 21].
-Therefore, we integrate a plain residual-in-residual block into the U-net cascade decoder in the PromptHSI to maintain the advantages of both designs without compromising.
-The versatility of the network can be ensured accordingly.
+Figure 3: Visualizations of the Fourier transforms of HSIs with specific degradations, displayed from top to bottom alongside the Fourier spectra of the residual images (obtained by subtracting the degraded images from the ground truth).
+3.1 Hybrid Architecture for HSI Restoration
+Effective HSI restoration requires an architecture capable of addressing diverse degradations within a single framework [8, 55].
+Existing approaches predominantly adopt (i) U-shaped encoder-decoder networks [54, 53], which effectively capture global semantic representations for overall restoration but have difficulty with fine details, or (ii) plain residual-in-residual networks [31, 19, 62], which enhance spatial features and high-frequency information but lack global awareness.
+AiO restoration models [47, 4, 15] and HSI-specific methods [39, 21, 24] favor the former, limiting fine-detail recovery.
+To overcome this, PromptHSI integrates residual-in-residual blocks within a U-shaped cascade decoder, preserving global context while refining local details for greater adaptability across degradation scenarios.
 3.2 Degradations Analysis in Frequency Domain
-Unlike RGB image restoration, HSI restoration requires careful consideration of spatial-spectral correlations across hundreds of bands.
-Our key insight comes from analyzing how different types of degradation manifest in the frequency-domain through Fourier analysis.
-Given a clean HSI $I_{\text{clean}}$, its Fourier transform can be expressed as:
-$F(I_{\text{clean}})=F_{s}(I_{\text{clean}})+iF_{c}(I_{\text{clean}})$ (1)
-where $F_{s}$ and $F_{c}$ represent the sine and cosine components respectively.
-Through empirical analysis, as shown in Figure 3, we observed that different types of degradation exhibit distinct patterns in how they affect these frequency components.
-Specifically, the relationship between degraded HSI $I_{\text{degraded}}$ and clean HSI $I_{\text{clean}}$ in frequency-domain can be characterized by an affine transformation:
-$F(I_{\text{degraded}})=\hat{\lambda}\odot F(I_{\text{clean}})+\hat{\mu}=(1+\lambda)\odot F(I_{\text{clean}})+\mu$ (2)
-where $\odot$ denotes element-wise multiplication, $\lambda$ and $\mu$ are degradation-specific parameters that: $\lambda$ represents the scaling factor that modulates frequency amplitudes, $\mu$ is the bias term that shifts the frequency distribution.
-Different types of degradation show characteristic patterns in $\lambda$ and $\mu$, such as ❶ Blur mainly affects high frequencies: $\lambda_{high}$ $\ll$ $\mu_{low}$.
-❷ Noise adds uniform perturbations: $\mu_{noise}$approx constant $c$.
-❸ Missing information creates structured patterns in both $\lambda$ and $\mu$.
+Unlike RGB image restoration, HSI restoration must account for complex spatial-spectral correlations across hundreds of spectral bands [58].
+The degradation process can be mathematically formulated as $I_{\text{degraded}}=\mathcal{D}(I_{\text{clean}})+\epsilon$, where $\mathcal{D}(\cdot)$ represents the degradation operator, encompassing effects such as blurring, stripe noise, and occlusion, while $\epsilon$ denotes additive noise.
+Frequency modulation.
+Recovering $I_{\text{degraded}}$ from an arbitrary degradation function $\mathcal{D}^{-1}\{\cdot\}$ is inherently challenging, particularly in the presence of composite degradations, where multiple degradation factors (e.g., noise, blur, and missing bands) interact in a complex manner.
+Unlike isolated degradations, which can often be addressed with predefined priors, composite degradations induce coupled distortions across both spatial and spectral dimensions, making traditional inverse modeling impractical.
+Our empirical analysis (refer to Figure 3) reveals that different types of degradation exhibit distinct but overlapping spectral characteristics.
+Motivated by this, we reformulate Equation 2 by modeling the degradation process as an affine transformation in the frequency domain:
+$\footnotesize F(I_{\text{degraded}})=\hat{\lambda}\odot F(I_{\text{clean}})+\hat{\mu}=(1+\lambda)\odot F(I_{\text{clean}})+\mu,$ (2)
+where $F(\cdot)$ denotes the Fourier transform, $\lambda$ modulates frequency amplitudes, $\mu$ shifts the frequency distribution, and $\odot$ represents element-wise multiplication.
+Under composite degradations, these frequency parameters exhibit complex dependencies:
+  • Blurring attenuates high-frequency components ($\lambda_{\text{blur}}\ll\mu_{\text{blur}}$), reducing fine details.
+  • Noise induces broadband perturbations, where ($\mu_{\text{noise}}\approx constant$) remains nearly uniform across frequencies.
+  • Occlusion and missing bands cause structured disruptions in both $\lambda$ and $\mu$, leading to spectral inconsistencies.
+  • Composite degradations result in frequency interactions, where overlapping effects create non-trivial distortions beyond individual degradations.
+For a clearer understanding, Figure 4 (left) visualizes these degradation-specific frequency characteristics, highlighting the non-additive nature of composite degradations in the Fourier domain.
+Figure 4: A diagram illustrating our proposed strategy, where integrating prompt guidance with frequency modulation facilitates the adaptive selection of $\lambda$ and $\mu$, enhancing HSI restoration under various degradation types.
+This mechanism encourages the model to efficiently converge toward optimal parameters.
+Figure 5: The architecture of the proposed PromptHSI, which is designed to tackle the composite degradations within a single universal model via VLM-guided frequency modulation.
+First, we design a network based on U-net-like encoder-decoder architecture for capturing global spatial-spectral information, while combining plain-residual-in residual networks in the proposed PGFAB for spatial-spectral enhancement.
+Afterwards, text-prompt-guidance and frequency-aware feature modulation is proposed to jointly better capture the representation of composite degradation with controllability.
+Prompt guidance.
+The diverse nature of HSI degradations presents significant challenges in learning the corresponding frequency coefficients.
+As shown in Figure 4 (right), previous AiO restoration methods [4, 11] utilize adaptive modules to classify degradation types.
+However, lacking explicit prompt-based guidance, these methods rely solely on implicit feature learning, making parameter selection unreliable, particularly under composite degradations, where multiple degradation effects exhibit complex spectral interactions.
+To address this, we introduce a VLM-driven prompt guidance mechanism, a core component of PromptHSI, which embeds frequency-aware priors into restoration.
+Leveraging textual cues from a given prompt $T$, our approach modulates frequency coefficients, refining the parameter space for faster convergence and enhanced frequency adaptation.
+This ensures precise control, adaptability, and robust performance across diverse degradations.
 4 Proposed Method
-Figure 4: The architecture of the proposed PromptHSI, which is designed to tackle the composite degradations within a single universal model.
-First, we design a network based on U-net-like encoder-decoder architecture [46, 45] for capturing global spatial-spectral information, while combining Residual-Dense Group (RDG) and Fast-Residual-Dense Block (FRDB) presented in DRCT [17] and RTCS [16], respectively, in the proposed PGFAB for spatial-spectral enhancement.
-Afterwards, text-prompt-guidance and frequency-aware feature map modulation is proposed to jointly better capture the representation of composite degradation with controllability.
 Overview.
@@
 Spatial and spectral feature enhancement.
-The output of PGFM is first improved by consecutive RDG [17] and FRDB [16] to improve fine-grained details.
-Then these features are interacted by cross-attention layer, which aims to comprehensively integrate the spatial-spectral feature information from different branches across all dimensions.
-Specifically, it can be expressed as follows:
-$\{Q_{\alpha},K_{\alpha},V_{\alpha}\}=\mathcal{F}^{\alpha}_{qkv}(F_{\alpha}),\quad\{Q_{\beta},K_{\beta},V_{\beta}\}=\mathcal{F}^{\beta}_{qkv}(F_{\beta}),$ (3)
-where $F_{\alpha},F_{\beta}$ denote features from the spatial and spectral enhancement block. Subsequently, we exchange the queries $Q$ of two branches for spatial-spectral interaction:
-$F_{f}^{\beta}=softmax(\frac{Q_{\alpha}K_{\beta}}{d_{k}})V_{\beta},\quad F_{f}^{\alpha}=softmax(\frac{Q_{\beta}K_{\alpha}}{d_{k}})V_{\alpha},$ (4)
-where $d_{k}$ is the scaling factor.
-Subsequently, we concatenate the results obtained by the cross-attention through $F_{f}^{\gamma}=Concat({F_{f}^{\beta},F_{f}^{\alpha}})$ to obtain the integrated features.
-$\{Q_{\gamma},K_{\gamma},V_{\gamma}\}=\mathcal{F}^{\gamma}_{qkv}(F_{f}^{\gamma}),\quad\hat{F_{f}^{\gamma}}=softmax(\frac{Q_{\gamma}K_{\gamma}}{d_{k}})V_{\gamma},$ (5)
-Finally, we employ spectral-spatial self-attention for feature co-enhancement to improve representation capability.
-Through a dual-branch architecture, we integrate text-prompt guidance effectively across both spatial and spectral dimensions.
-By incorporating plain-residual-in-residual [27, 17, 16] modules within the U-shape cascade decoder [46, 8], we enhance fine-grained image restoration while integrating global information, thereby effectively addressing various degradations for HSI restoration.
+The PGFM output is first refined using sequential RDG [19] and FRDB [18] modules to enhance fine details.
+Next, a cross-attention layer integrates spatial and spectral features from different branches.
+Specifically, we compute:
+$\footnotesize\{Q_{\alpha},K_{\alpha},V_{\alpha}\}=\mathcal{F}^{\alpha}_{qkv}(F_{\alpha}),\quad\{Q_{\beta},K_{\beta},V_{\beta}\}=\mathcal{F}^{\beta}_{qkv}(F_{\beta}),$ (3)
+where $F_{\alpha}$ and $F_{\beta}$ are the spatial and spectral features, respectively. We then exchange queries between branches:
+$\footnotesize F_{f}^{\beta}=\operatorname{softmax}\left(\frac{Q_{\alpha}K_{\beta}}{d_{k}}\right)V_{\beta},\quad F_{f}^{\alpha}=\operatorname{softmax}\left(\frac{Q_{\beta}K_{\alpha}}{d_{k}}\right)V_{\alpha},$ (4)
+where $d_{k}$ is a scaling factor. The outputs $F_{f}^{\alpha}$ and $F_{f}^{\beta}$ are combined into $F_{f}^{\gamma}$, which is then enhanced through cross-attention for feature intercalation:
+$\footnotesize\{Q_{\gamma},K_{\gamma},V_{\gamma}\}=\mathcal{F}^{\gamma}_{qkv}(F_{f}^{\gamma}),\quad\hat{F}_{f}^{\gamma}=\operatorname{softmax}\left(\frac{Q_{\gamma}K_{\gamma}}{d_{k}}\right)V_{\gamma}.$ (5)
+Our dual-branch architecture effectively integrates text-prompt guidance across spatial and spectral dimensions.
+By embedding plain-residual-in-residual modules [31, 19, 18] within a U-shaped cascade decoder [54, 8], our framework achieves both fine-grained and global restoration, effectively addressing various HSI degradations.
 Table 1: Overall performance evaluation for composite degradation and complexity comparison.
@@
 In this study, we employ CLIP [48], which integrates vision and language information and performs effectively in text feature extraction.
-We freeze the pre-trained CLIP to maintain great linguistic consistency and add additional MLP layer $\Phi_{t}^{adpt.}$ as the feature adapter for pattern matching.
+We freeze the pre-trained CLIP text encoder $\mathcal{E}^{CLIP}_{text}$ to maintain great linguistic consistency and add additional MLP layer $\Phi_{t}^{adpt.}$ as the feature adapter for pattern matching.
 With {}$\{\cdot\}_{e}$ denoting the frozen weights, this process can be expressed as:
-$\footnotesize F_{text}=\Phi_{t}^{adpt.}(\{\mathcal{F}^{CLIP}_{text}\}_{e}(T_{text})),$ (6)
+$\footnotesize F_{text}=\Phi_{t}^{adpt.}(\{\mathcal{E}^{CLIP}_{text}\}_{e}(T_{text})),$ (6)
 where $F_{text}$ denotes the task descriptors. With the task descriptors available, the next challenge is how to effectively incorporate this guidance into the HSI restoration process.
@@
 Given an input feature map $F_{f}$, we first decompose it into frequency components:
-$F_{f}=F_{f}^{l}+F_{f}^{h},$ (7)
+$\footnotesize F_{f}=F_{f}^{l}+F_{f}^{h},$ (7)
 where $F_{f}^{l}$ and $F_{f}^{h}$ represent low and high frequency components, obtained through frequency-domain filtering.
-Afterward, we design the MLP layer $\Psi_{m}^{i}$ to extract this connection with the intensity and bias embeddings for feature modulation. Therefore, it can be obtained:
-$\lambda_{m}=\Psi_{m}^{I}(F_{text}),\ \mu_{m}=\Psi_{m}^{II}(F_{text}),$ (8)
+Afterward, to merge text and image features, previous work [57] proposed employing a single MLP layer, $\Psi_{m}^{i}$, to integrate them with intensity and bias embeddings.
+Consequently, our feature modulation can be represented as:
+$\footnotesize\lambda_{m}=\Psi_{m}^{I}(F_{text}),\ \mu_{m}=\Psi_{m}^{II}(F_{text}),$ (8)
 where $\Psi_{m}^{I}$ and $\Psi_{m}^{II}$ are the chunk operations of $\Psi_{m}$ to form the intensity and bias controller embeddings.
-To address the spectral consistency challenge in HSIs, our frequency-aware feature modulation adopts an affine transformation model tailored for spectral data, formulated as:
-$\hat{F}_{f}=(1+\lambda_{m}^{l})\odot F_{f}^{l}+(1+\lambda_{m}^{h})\odot F_{f}^{h}+\mu,$ (9)
+To address the spectral consistency challenge in HSIs, our frequency-aware feature modulation adopts an affine transformation model (Eq.
+2) tailored for spectral data, formulated as:
+$\footnotesize\hat{F}_{f}=(1+\lambda_{m}^{l})\odot F_{f}^{l}+(1+\lambda_{m}^{h})\odot F_{f}^{h}+\mu,$ (9)
 where Intensity controller $\lambda$in [$\lambda_{m}^{l}$, $\lambda_{m}^{h}$] modulates the amplitude of high-frequency components to address noise, while Bias controller $\mu$ shifts the low-frequency baseline to compensate for blurring and band-missing effects.
@@
 4.3 Loss Function
-Given the wide range of spectra inherent in HSI, we utilize only the $\ell_{1}$ loss for training.
-However, this strategy may not adequately capture the variability across different bands.
-Consequently, the $\ell_{1}$ loss might inadvertently focus on minimizing bands with larger values, affecting the spectral integrity of specific bands.
-Spectral angle mapper loss.
-SAM metric is also critical for material identification and less sensitive to intensity variations.
-The $\ell_{\text{SAM}}$ loss prioritizes angular discrepancies over magnitudes, offering a reliable metric to evaluate the spectral fidelity of the reconstructed HSI.
-$\ell_{\text{SAM}}=\frac{1}{N}\sum_{n=1}^{N}\cos^{-1}\left(\frac{I_{n}^{T}I^{*}_{n}+\epsilon}{\|I_{n}\|_{2}\cdot\|I_{n}^{*}\|_{2}+\epsilon}\right),$ (10)
-where $I_{n}$ and $I^{*}_{n}$ represent the $n$-th spectral vector of the ground-truth and the reconstructed HSI, respectively. The term $\epsilon$ is included to avoid divergence.
-Stationary wavelet transform loss.
-Wavelet sub-band alignment is able to manage various aspects of information encoded in HSIs.
-On the other words, it facilitates more detailed supervision in the frequency-domain.
-$\ell_{\text{SWT}}=\sum_{j}\lambda_{j}\big{\|}\text{SWT}(I)_{j}-\text{SWT}(I^{*})_{j}\big{\|}_{1},$ (11)
-where $\lambda_{j}$ represents the count of sub-bands decomposed via stationary wavelet transformation.
-Band-wise MSE loss.
-The band-wise disparity helps stabilize the network optimization process and is advantageous to achieve convergence, balancing restoration quality across all spectra.
-The Band-wise MSE loss can be simply defined as follows:
-$\ell_{\text{BMSE}}=\frac{1}{N}\sum_{n=1}^{N}{\|I_{n}-I_{n}^{*}\|_{2}},$ (12)
-Total loss. Ultimately, the total loss function for training the proposed PromptHSI $\ell_{\text{total}}$ can be expressed as:
-$\ell_{\text{total}}=\lambda_{1}\ell_{\text{1}}+\lambda_{2}\ell_{\text{SAM}}+\lambda_{3}\ell_{\text{SWT}}+\lambda_{4}\ell_{\text{BMSE}},$ (13)
-where $\lambda_{1,2,3,4}$ serve as the weighting coefficients to equilibrate $\ell_{1}$, $\ell_{\text{SAM}}$, $\ell_{\text{SWT}}$, and $\ell_{\text{BMSE}}$, set at 1, 0.1, 0.01, 0.01 throughout all experiments, respectively.
+Given the wide spectral range in HSIs, using only the $\ell_{1}$ loss during training may not fully capture band variability; it tends to prioritize bands with larger values, potentially compromising spectral integrity.
+Spectral Angle Mapper Loss. The SAM metric is essential for material identification and is less sensitive to intensity variations. It emphasizes angular discrepancies over magnitudes:
+$\footnotesize\ell_{\text{SAM}}=\frac{1}{N}\sum_{n=1}^{N}\cos^{-1}\left(\frac{I_{n}^{T}I^{*}_{n}+\epsilon}{\|I_{n}\|_{2}\cdot\|I_{n}^{*}\|_{2}+\epsilon}\right),$ (10)
+where $I_{n}$ and $I^{*}_{n}$ denote the $n$-th spectral vectors of the ground-truth and reconstructed HSIs, respectively, and $\epsilon$ prevents divergence.
+Stationary Wavelet Transform Loss. SWT loss enforces frequency-domain consistency via wavelet sub-band alignment:
+$\footnotesize\ell_{\text{SWT}}=\sum_{j}\lambda_{j}\big{\|}\text{SWT}(I)_{j}-\text{SWT}(I^{*})_{j}\big{\|}_{1},$ (11)
+with $\lambda_{j}$ weighting the contribution of each sub-band.
+Band-wise MSE Loss. BMSE loss balances reconstruction across all spectral bands to stabilize optimization:
+$\footnotesize\ell_{\text{BMSE}}=\frac{1}{N}\sum_{n=1}^{N}\|I_{n}-I_{n}^{*}\|_{2}.$ (12)
+Total Loss.
+The overall loss for training PromptHSI is defined as $\ell_{\text{total}}=\lambda_{1}\ell_{\text{1}}+\lambda_{2}\ell_{\text{SAM}}+\lambda_{3}\ell_{\text{SWT}}+\lambda_{4}\ell_{\text{BMSE}}$, where the coefficients are set to $\lambda_{1}=1$, $\lambda_{2}=0.001$, $\lambda_{3}=0.01$, and $\lambda_{4}=0.01$ throughout all experiments.
 5 Experimental Results
@@
 This study employed the AVIRIS dataset provided by Lin et al.
-[30, 18], an extensive HSI dataset, with numerous bands spanning diverse geographical areas for experiments.
-It includes 2,078 HSIs, which were randomly divided into training, validation, and testing sets.
-The training set consisted of 1,678 images and both the validation and testing sets contained 200 images each.
-As illustrated in Supplementary, all clean HSI was pre-processed to synthesize the composite degraded HSIs and corresponding text descriptions.
+[35, 20], an extensive HSI dataset with numerous bands spanning diverse geographical areas, for experiments.
+The dataset comprises 2,078 HSIs, which were randomly divided into training (1,678 images), validation (200 images), and testing (200 images) sets.
+As illustrated in the Supplementary, all clean HSIs were pre-processed to synthesize composite degraded HSIs along with corresponding text descriptions.
+This dataset is publicly available to facilitate further research.
 5.1 Quantitative Results and Analysis
 To evaluate the performance of PromptHSI, we compared it with eight other supervised HSI restoration methods for specific degradation, including HSIDwRD [60], SERT [29], SRDNet [39], SNLSR [21], SGNet [43], AACNet [56], SST [28], and HSDT [26] in the synthesized AVIRIS dataset with composite degradations.
-Because there is no AiO HSI restoration method, we compare four AiO RGB image restoration methods in the experiments: AirNet [23], PromptIR [39], HAIR [5], InstructIR [10], and OneRestore [13].
+Because there is no AiO HSI restoration method, we compare four AiO RGB image restoration methods in the experiments: AirNet [27], PromptIR [47], HAIR [4], InstructIR [10], AdaIR [11], and OneRestore [15].
 The model performance was measured using the four metrics, including PSNR, SAM, RMSE, and ERGAS.
@@
 Therefore, these models are not effective in transferring to AiO HSI restoration.
-Specifically, OneRestore [13] incorporated task description features into its encoder block, which hindered its capacity to accurately capture the spatial-spectral correlation in HSI because of the domain gap with HSI features, resulting in decoding failures.
-Moreover, it lacks considerations for tasks requiring high-frequency detail reconstruction, potentially limiting its effectiveness in the spatial or spectral super-resolution task.
+Specifically, OneRestore [15] integrated task description features into its encoder, which hindered its ability to capture spatial-spectral correlations in HSIs due to a domain gap, leading to decoding failures.
 5.2 Ablation Studies and Analysis
@@
 We further conduct ablation studies to investigate the impact of different components and strategies.
-Integration strategy of feature-map modulation.
-As shown in Table 4, our proposed frequency-aware feature modulation demonstrates superior effectiveness compared to the weighted-sum manner, where we implemented an MLP layer for dynamic weighted averaging of inputs and outputs.
+Integration strategy of frequency-aware modulation.
+As shown in Table 12, modulating frequency via prompting demonstrates superior effectiveness compared to the weighted-sum manner, where we implemented an MLP layer for dynamic weighted averaging of inputs and outputs.
 The intensity-only configuration outperforms the bias-only, with a notably larger advantage in the ERGAS metric, suggesting that this integration strategy effectively guides network restoration across most scenarios.
 Visual-prompt or text-prompt.
-The type of prompting is crucial, as it determines its ability to recognize various degradations and effectively guide and control model behavior.
-In the first configuration, we select RGB bands from HSIs as visual prompts, encoding them through CLIP [40] and using the proposed PGFM to guide the model.
-Secondly, we encode both RGB bands and textual prompts to generate hybrid task descriptors for guidance.
-In the final configuration, we remove PGFM and use the PIM and PGM block proposed in PromptIR [39] for the generation of visual prompts and interaction of features.
-The experimental results are presented in Table 4.
-Due to multiple or composite degradations in the spatial-domain leading to serious information loss, it became difficult to obtain compact task descriptors or valid visual prompts, ultimately resulting in diminished prompt effectiveness and compromised restoration.
-The use of textual features helps to circumvent these issues and achieves robust prompting guidance through semantic information.
+The type of prompting is crucial for recognizing degradations and guiding model behavior.
+In our first configuration, we use RGB bands from HSIs as visual prompts, encoded via CLIP [48] and guided by PGFM.
+In the second, we combine RGB and textual prompts to form hybrid task descriptors.
+In the final configuration, PGFM is removed and we employ the PIM and PGM blocks from PromptIR [47] for generating visual prompts and feature interaction.
+Table 4 shows that composite spatial degradations cause significant information loss, making it difficult to obtain compact task descriptors or effective visual prompts.
+Incorporating textual features overcomes these challenges by providing robust semantic guidance.
 Ablation of network component.
@@
 6 Conclusion
-In this paper, we present PromptHSI, the first universal framework for hyperspectral image restoration that effectively addresses composite degradations, which conventional HSI restoration methods have failed to handle.
-Unlike existing All-in-One RGB image restoration approaches that struggle with domain gaps between RGB and HSI features and structures, our method is specifically designed to address HSI-specific attributes and commonly encountered degradation patterns.
-Then, our method is controllable through text-prompt guidance, based on a novel frequency-aware modulation strategy that effectively bridges the gap between text and HSI through frequency pattern matching and distribution adjustment in the frequency-domain.
-Our approach demonstrates superior performance across various types of degradation and their combinations, significantly outperforming existing methods.
+In this paper, we introduced PromptHSI, the first universal framework for hyperspectral image restoration that effectively addresses composite degradations—a challenge that traditional methods have struggled to overcome.
+Unlike existing AiO RGB restoration approaches, which are hampered by domain gaps and inadequately capture HSI-specific degradation patterns, our method is specifically tailored to the unique attributes of HSIs.
+By integrating frequency-aware modulation with vision-language model–guided prompt learning, PromptHSI bridges textual cues and frequency-domain information, enabling adaptive and controllable restoration.
+Our approach significantly outperforms state-of-the-art methods across various degradation scenarios.
+Moreover, we present and release the first composite degradation dataset for HSI restoration, paving the way for future research in this field.
+Supplementary Materials
+Figure 10: The t-SNE map visualization for the proposed PromptHSI and other All-in-One RGR image restoration methods.
+Note that for training PromptHSI, the generated dataset includes various degradations, with each type presented separately instead of concurrently.
+Subsequently, we visualize the t-SNE map with ❶ Top: multiple degradation; ❷ Bottom: composite degradation during testing-time.
+Appendix A Appendix
+Overview.
+This supplementary material provides a more detailed explanation of the model configuration, the preparation of the dataset, and additional results of the experiments.
+❶ First, we provide a detailed experiment setup, dataset description, and its preparation pipeline in Sections B and C.
+❷ Secondly, we provide the illustration of the PromptHSI decoder module and present our model configuration in Section D.
+❸ Next, we present additional experiment results and visualization in Section E.
+Appendix B Experiment Setup
+Dataset description.
+This study used a dataset captured by the Airborne Visible/Infrared Imaging Spectrometer AVIRIS Data Portal: https://aviris.jpl.nasa.gov/data/index.html(AVIRIS) [52, 35].
+The HyperAD: https://github.com/chingheng0808/HyperADdataset encompasses a wide range of natural settings throughout the United States and Canada, including urban areas, mountains, lakes, agricultural regions, and diverse vegetation, gathered from 2006 to 2011.
+The original HSIs were partitioned into distinct subimages, each measuring 259 $\times$ 259 pixels, composed of 224 spectral bands spanning wavelengths from 400 to 2500 nm.
+Consistent with prior research [34, 35, 20], spectral bands identified as lower quality (1-10, 104-116, 152-170, and 215-224) were excluded, resulting in HSIs with 172 spectral bands.
+Hyperparameter setting.
+We used single NVIDIA-A100 for training.
+The batch size was set to 8, and the number of training epochs was fixed at 300 for all experiments involving the proposed method.
+The AdamW optimizer [41] ($\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay $1\times 10^{-2}$) was employed, with an initial learning rate of $3\times 10^{-4}$.
+The learning rate was dynamically adjusted during the training process using a step-based learning rate scheduler.
+The scheduler reduces the learning rate by a factor of 0.5 every 50 epochs to ensure stable convergence.
+Data augmentation.
+The input HSI was randomly cropped to patches with the size of $112\times 112\times 172$, incorporating random horizontal, vertical rotations and flipping during the training phase.
+During inference, the model was evaluated using input dimensions of $224\times 224\times 172$.
+Table 5: Inference time comparison.
+For our simulation, we use randomly generated tensor with the size of $224\times 224\times 172$, conducting inference $30$ times and taking the average for evaluation on All-in-One RGR image restoration methods.
+Method Inference time per inputs ↓ ($\mu$ $\sigma$)
+AirNet [27] 230.54 ± 59.61 ms
+PromptIR [47] 178.58 ± 47.92 ms
+HAIR [4] 162.33 ± 38.88 ms
+AdaIR [11] 107.93 ± 59.86 ms
+InstructIR [10] 32.62 ± 8.63 ms
+OneRestore [15] 24.90 ± 5.94 ms
+PromptHSI (Ours) 282.78 ± 74.48 ms
+Appendix C Dataset Preparation
+Figure 11: The proposed dataset synthesis pipeline.
+Each degradation is activated with probability $p$.
+Note that we set $p$ = $0.5$ for all experiments.
+For the diversity of synthetic dataset, each active degradation chooses and implements one of its subdegradations on the input HSI.
+To the best of our knowledge, there have been no reliable simulations and analysis of composite degradation in HSI restoration.
+Therefore, designing a realistic degradation process that accurately simulates real-world scenarios is important.
+To address this, we modify the GDM (Gated Degradation Model) [61] from the RGB image blind super-resolution task to accommodate our task.
+C.1 Degradation Model
+It is essential to consider the complete imaging acquisition pipeline when simulating the hyperspectral imagery process.
+Specifically, the observed images are initially affected by cloud coverage or shadow occlusion, resulting in ground surface invisibility.
+Subsequently, the image resolution is altered due to hardware limitations.
+The imaging components then introduce a random noise into the image.
+Finally, at the transmission, atmospheric interference or bandwidth constraints lead to the manifestation of stripe noise.
+Classical hyperspectral imaging model.
+HSI restoration is an ill-posed problem, which assumes the clean image is affected by multiple types of degradation.
+Mathematically, the degraded image $I_{\text{degraded}}$ is generated from the clean image $I_{\text{clean}}$ as follows:
+$\footnotesize I_{\text{degraded}}=D_{stripe,n,s,c}(I_{\text{clean}})=\left[\downarrow(c\oplus I_{\text{clean}})+n\right]\odot M_{stripe},$ (13)
+First, clean HSI $I_{\text{clean}}$ is degraded by cloud occlusion (denoted by $c$).
+Then, the occluded HSI is down-sampled (denoted as ($\downarrow$($\cdot$)↓ ( ⋅ )) and an additive white Gaussian noise (denoted as $n$) is added to the degraded HSI.
+Finally, some of the HSI $I_{\text{degraded}}$ bands are corrupted and missed by the stripe effect during data transmission (denoted $stripe$).
+Cloud occlusion degradations.
+We followed the Satellite Cloud Generator [12] to synthesize the cloud and degradation.
+The cloudy HSI $I_{\text{cloudy}}$ is generated by blending the clear HSI $I_{\text{clean}}$ with the cloud HSI $I_{\text{cloud}}$ using the cloud transparency mask $M_{C}$:
+$\footnotesize I_{\text{cloudy}}=(1-M_{C})\cdot I_{\text{clean}}\oslash M_{C}\cdot I_{\text{cloud}},$ (14)
+[…]
+Noise and stripe effect degradations.
+Noise introduced during data acquisition can significantly affect image quality.
+Two common types of noise in HSI are Additive White Gaussian Noise (AWGN) and stripe noise.
+Stripe pattern noise appears as vertical or horizontal lines in specific bands, typically caused by sensor calibration problems and data transmission.
+When both AWGN and stripe pattern noise are present in spatial location $(i,j)$ in spectral band $k$, the noisy signal can be represented as:
+$\footnotesize I_{\text{noisy}}(i,j,k)=\left[I_{\text{clean}}(i,j,k)+N_{\text{AWGN}}(i,j,k)\right]\odot M_{S}.$ (15)
+Probabilistic gated degradation model.
+Following GDM [61], which aims to synthesize the dataset with composite degradations for RGB image, we redesign its degradation types to better suit the attribute of HSI.
+The gate mechanism is introduced to randomly select the base degradation types to be included in the degradation process.
+Formally, the probabilistic gated degradation process $D_{G}$ can be expressed as:
+$\footnotesize I_{\text{degraded}}=D_{G}(I_{\text{clean}})=(\sigma_{g}(D_{1})\circ\sigma_{g}(D_{2})\cdots\sigma_{g}(D_{m}))(I_{\text{clean}}),$ (16)
+where $\sigma_{g}(D_{i})(r^{d})=\begin{cases}D_{i}(r^{d}),&g_{i}\sim\text{Bernoulli}(p),\\ r^{d},&g_{i}=0,\end{cases}$ and $r^{d}$ denotes the input HSI.
+In this case, each gate $g_{i}$ is activated with a probability $p$ (e.g., $p=0.5$), which determines whether a specific degradation $D_{i}$ is applied.
+PGDM enables the diverse integration of fundamental degradation types, enhancing the synthesized HSI to be more vibrant in real-world scenarios.
+Table 6: The details configurations and hyperparameter settings for degradation synthesis.
+The configuration for the Cloudy degradation is based on the parameters of Satellite Cloud Generator [12].
+As for $K$ in the band-missing configuration, it represents a value between 0 and 172 (the total number of HSI bands).
+The last column displays the PSNR and SAM metrics for the degraded HSIs, which contain only the corresponding degradation in the testing set.
+  Degradations Types Configurations Mean PSNR / SAM on specific degradation Cloudy Thickly locality_degree=[2, 4], min_lvl=0.0, max_lvl=1.0, clear_threshold=[0.0, 0.4], blur_scaling=1, cloud_color=True, channel_offset=0, decay_factor=1.0 11.0379 / 4.0143 Thinly locality_degree=1, min_lvl=[0.0, 0.4], max_lvl=[0.4, 0.6], clear_threshold=0.0, blur_scaling=2, cloud_color=True, channel_offset=0, decay_factor=1.0 7.0316 / 17.1393 Blurring Spatial Bilinear interpolation is used to downsample by a factor of 4, followed by bilinear interpolation to restore the image to its original size in the spatial domain.
+22.2780 / 3.8414 Spectral Using a Gaussian window filter with a window size of 5 and stride of 4 to smooth along the spectral dimension, followed by nearest-neighbor interpolation to restore the image to its original size in the spectral domain.
+24.8798 / 12.6105 Noisy Gaussian noise Using SNR to control the noise intensity, where $\text{SNR}\sim\mathcal{N}(\mu,\sigma)$, with $\mu=35$ and $\sigma=5$.
+16.2706 / 12.9214 Band-missing Complete All pixel values in randomly chosen $K$ bands are set to 0.
+28.2688 / 26.6282 Band-wise On randomly chosen $K$ bands, pixels in a predetermined sequence of rows (for instance, the 1st, 3rd, 5th, 7th, 9th, etc.) are set to 0.
+30.1465 / 5.5401 Partial On randomly chosen $K$ bands, each row in the bands has a 0.3 probability of being set to 0.
+27.8382 / 5.8436
+C.2 Degradation Simulation
+Following the exposition in the previous subsection, where we detailed the simulation methodology for hyperspectral image degradation, we now proceed to elaborate on the considerations and processes involved in simulating each specific type of degradation.
+C.2.1 Cloud Occlusion
+Background.
+Although optical satellite imagery can be applied to various tasks, it is often affected by weather conditions, such as cloud cover, making it difficult to clearly observe the appearance of surface features.
+As a result, compiling high-quality datasets that are affected by cloud cover proves to be a difficult task, making the generation of high-quality cloudy data necessary.
+Synthesis.
+Following the method for synthesizing clouds in The official Github repository of Satellite Cloud Generator [12]: https://github.com/strath-ai/SatelliteCloudGenerator.git Satellite Cloud Generator [12], we first use a pair of real HSIs, one containing clouds and the other a neighboring image.
+Afterwards, using the cloud mask detector proposed in [51], we calculate the cloud mask area in the spectral bands of Sentinel-2.
+This mask is then used to approximate the cloud end-member function in different spectra, and finally, we generate clouds on the given clean HSI that contain thinly or thickly cloudy HSI, as shown in Figure 12.
+Figure 12: Visualization of synthetic cloudy HSI.
+The first two rows illustrate thick cloud cover conditions, whereas the last two rows demonstrate thin cloud cover conditions.
+Thick clouds cover a smaller portion of the overall HSI, but the areas obscured by these clouds are difficult to observe.
+In contrast, thin clouds, while covering a larger area, allow for faint visibility of the underlying surface features, somewhat resembling a foggy effect.
+Figure 13: Visualization of simulated band missing on HSI.
+We selected multiple bands to demonstrate different types of band missing degradation, with the 6th, 15th, and 25th bands representing the B, G, and R bands, respectively.The first two rows depict the band complete missing case, the next two rows indicate instances of band-wise missing, and the final two rows represent cases of partial missing.
+C.2.2 Spatial Blurring and Spectral Blurring
+Background.
+Due to diverse sensor payloads and varying orbital altitudes of Low Earth Orbit (LEO) satellites, remote sensing data exhibits significant heterogeneity in both spatial and spectral resolutions.
+Hyperspectral imaging sensors, in particular, often suffer from limited spatial resolution due to technical and cost constraints, hindering detailed surface feature observation.
+To address these limitations, we incorporate both spatial super-resolution and spectral super-resolution tasks into our model framework.
+Synthesis.
+To simulate resolution degradation, we implement a downsampling-upsampling procedure across spatial or spectral domains.
+First, the formulations for spatial blurring process can be defined as follows:
+$\footnotesize I_{\text{spatial blurring}}={I_{\text{clean}}}_{Spa_{\downarrow}^{4}Spa_{\uparrow}^{4}},$ (17)
+where $Spa_{\downarrow}$ and $Spa_{\uparrow}$ denotes as spatial bilinear down-sampling and bilinear interpolation up-sampling with factor of $4$, respectively.
+Subsequently, the spectral blurring process can be expressed as:
+$\footnotesize I_{\text{spectral blurring}}=(\Phi(I_{\text{clean}},5,4))_{Spe_{\uparrow}^{4}},$ (18)
+where $\Phi(\cdot,\omega,\varsigma)$ denotes a sliding Gaussian window filter with window size $\omega$ and stride $\varsigma$, respectively, to reduce the spectral resolution.
+$Spe_{\uparrow}$ represents nearest-neighbor interpolation with factor of $4$ along the spectral dimension to restore the original spectral size.
+C.2.3 Random Noise
+Background.
+In hyperspectral imaging, captured HSI often contains noise due to environmental interference, lower-quality sensors, or signal transmission issues.
+Thus, HSI denoising is crucial for HSI restoration and its applications.
+Synthesis.
+To enhance the diversity of noise magnitude in our dataset, we simulate AWGN using Signal-to-Noise Ratios (SNR) sampled from a normal distribution with mean $\mu$ and standard deviation $\sigma$ to control the noise intensity.
+The Gaussian noise $I_{\epsilon}$, randomly drawn from $\mathcal{N}(\mathbf{0},\mathbf{1})$, is scaled by the intensity term to generate the AWGN.
+The corresponding formula and its effect on the HSI are expressed as:
+$\footnotesize I_{\text{noisy}}=I_{\text{clean}}+N_{\text{awgn}}=I_{\text{clean}}+I_{\epsilon}\cdot\sqrt{\frac{\|I_{\text{clean}}\|_{2}^{2}}{\text{SNR}}},$ (19)
+C.2.4 Band-Missing
+Background. Due to the conditions of satellite sensors and signal interference during transmission, HSIs often encounter band-missing issues in addition to noise [59].
+Synthesis.
+We simulate this degradation and categorize it into three types as follows [18].
+The first is where the pixel values of certain bands are completely missing, referred to as ❶ Band Complete Missing.
+The second and third types involve missing pixel values in certain rows of the image, resulting in stripe-like defects.
+These are classified as ❷ Band-Wise Missing and ❸ Partial Missing, based on whether the rows disappear in a regular or irregular pattern.
+The visualization of the three different missing types are shown in Figure 13.
+Table 7: Text prompt formats.
+Different colors represent various types of degradation.
+If a particular degradation type is absent, it will be excluded from the text prompt.
+The notation [] specifies the subtype of the respective degradation type, while $N_{mb}$ denotes the number of bands affected by band-missing degradation.
+Prompt type Prompt format
+Long prompt ”This hyperspectral image faces with ’[band-missing type]’ on $N_{mb}$ bands; it also confronts ’[cloudy type], ’[noisy]’; besides, there exists ’blurring effect in [blurring type] domain’.
+”
+Short prompt [cloudy type], [noisy], [blurring type], [band-missing type]
+Table 8: Effect of text-prompt length.
+Length PSNR / SAM / RMSE / ERGAS #Params FLOPs
+Long 25.879 / 6.684 / 0.0196 / 10.789 26.15M 135.83G
+Short 26.443 / 6.060 / 0.0186 / 8.357 26.15M 135.83G
+Appendix D Network Details
+Figure 14: The detailed Architecture of our proposed Prompt-guided Feature Aggregation Block (PGFAB).
+In terms of our PromptHSI architecture, all configurations and output sizes of each layer are displayed in Table 9.
+Afterwards, we introduce the details of the Prompt-guided Feature Aggregation Block (PGFAB), as showan in Figure 14, which consists of the Prompt-guided Feature Modulation (PGFM) module and feature enhancement.
+The motivation of PGFM have been illustrated in our manuscript; here we further discuss the details of feature enhancement.
+Dual-branch design for feature aggregation.
+Inspired by SPACE [32], which uses two branches designed for efficient HSI compressed sensing.
+This design allows the network to focus independently on learning spectral and spatial features.
+We adopt the Residual Dense Group (RDG) [19] for the spatial feature extraction branch , and the Fast Residual Dense Block (FRDB) [18] for the spectral feature extraction branch.
+Residual dense group.
+This architecture incorporates Swin Transformer [31, 40] with dense connections, demonstrating a more lightweight and efficient parameter-size compared to the Residual Hybrid Attention Groups (RHAG) in HAT [7].
+This design exhibits superior capability in capturing global spatial details.
+Fast residual dense block.
+Built on group convolutions integrated with dense connections [22], excels at capturing local spatial-spectral details.
+In particular, the group convolution effectively leverages the low-rank prior [50, 25] inherent in HSI while achieving a reduced parameter-size compared to standard convolutions.
+It aligns with the HSI characteristic that spectral information exhibits high correlation in local regions while containing substantial redundancy in long-range interactions.
+Table 9: The configurations of the PromptHSI architecture.
+For some notations used in the Table, $H$, $W$, and $C$ represent the height, width, and the number of bands of the input HSI, respectively.
+$N_{\alpha}$ and $N_{\beta}$ denote the depth of the RDG [19] and FRDB [18] blocks in the dual-branch designed feature aggregation.
+$c$, $k$, $s$, $head$, $ws$, $ps$, and $n$ express the number of output channels, kernel size, stride, attention heads, window size, patch size, and the number of blocks, respectively.
+$ds$-$conv$ refers to depthwise separable convolution [9], $bn$ stands for batch normalization.For the task descriptor generation input, the text-prompt for degradation description follows a short format like ”$description^{1},\ldots,description^{n}$”, as shown in Table 8.
+  Models Layers Configurations Output size Input1 Degradation text prompt 1 string Task descriptor Text tokenizer BPE-based subword text tokenizer 77 generation CLIP text encoder [48] ViT-B/32 $512$ Adapter MLP1$(c=64)$, $leakyrelu$, MLP2$(c=512)$ $512$ PromptHSI Encoder Input2 Degraded hyperspectral image ${H}\times{W}\times{C}$ Conv1 $c=64,k=5,s=1$ ${H}\times{W}\times{64}$ SSAFEB1 $ds$-$conv(c=128,k=5,s=2),bn,leakyrelu$ ${\frac{H}{2}}\times{\frac{W}{2}}\times{128}$ SSAFEB2 $ds$-$conv(c=256,k=3,s=2),bn,leakyrelu$ ${\frac{H}{4}}\times{\frac{W}{4}}\times{256}$ SSAFEB3 $ds$-$conv(c=512,k=3,s=2),bn,leakyrelu$ ${\frac{H}{8}}\times{\frac{W}{8}}\times{512}$ Decoder PGFAB1 $c=256,head=8,ws=7,ps=4,N_{\alpha}=2,N_{\beta}=1$ ${\frac{H}{4}}\times{\frac{W}{4}}\times{256}$ Concatenation1 PGFAB1 + SSAFEB2 ${\frac{H}{4}}\times{\frac{W}{4}}\times{512}$ Conv2 $c=256,k=1,s=1$ ${\frac{H}{4}}\times{\frac{W}{4}}\times{256}$ PGFAB2 $c=128,head=8,ws=7,ps=4,N_{\alpha}=2,N_{\beta}=1$ ${\frac{H}{2}}\times{\frac{W}{2}}\times{128}$ Concatenation2 PGFAB2 + SSAFEB1 ${\frac{H}{2}}\times{\frac{W}{2}}\times{256}$ Conv3 $c=128,k=1,s=1$ ${\frac{H}{2}}\times{\frac{W}{2}}\times{128}$ PGFAB3 $c=D,head=4,ws=7,ps=4,N_{\alpha}=2,N_{\beta}=1$ ${H}\times{W}\times{64}$ Concatenation3 PGFAB3 + Conv1 ${H}\times{W}\times{128}$ Conv4 $c=128,k=1,s=1$ ${H}\times{W}\times{128}$ Transformer blocks $c=128,head=4,n=2$ ${H}\times{W}\times{128}$ Conv5 $c=C,k=1,s=1$ ${H}\times{W}\times{C}$
+Appendix E Additional Experiments
+In addition to the experimental details and ablation studies presented in the manuscript, this section conducts further experiments to evaluate ❶ the effects of different encoders, ❷ effect of the length of degradation descriptions, and ❸ effect of the feature enhancement.
+Rows are colored to distinguish different approaches: Adopted, Not adopted.
+E.1 Ablation Study
+We provide additional ablation studies and comparisons with various modules in Table 12, Table 12, and Table 12 to validate the effectiveness of these modules.
+Effect of text-prompt length.
+We further compared the impact of long and short text prompts for the model’s performance, where the formats of text-prompts and results are shown in Tables 7 and 8.
+The long text-prompt provides distinct descriptions for different degradation types, combining them into a single statement with a tone closer to natural human linguistic speech compared to the short prompt.
+Table 10: Impact of the type of convolution in SSAFEB. Note that S indicate standard convolution; DS represent depthwise separable convolution [9].
+Type PSNR / SAM / RMSE / ERGAS #Params FLOPs
+S 26.112 / 6.307 / 0.0186 / 10.258 27.65M 153.17G
+DS 26.443 / 6.060 / 0.0187 / 8.357 26.15M 135.83G
+Table 11: Ablation of feature co-enhancement.
+Spa Spe PSNR / SAM / RMSE / ERGAS #Params FLOPs
+$\times$ $\times$ 25.846 / 6.403 / 0.0191 / 11.514 22.70M 129.67G
+$\checkmark$ $\times$ 26.170 / 6.295 / 0.0187 / 12.911 24.08M 132.14G
+$\times$ $\checkmark$ 25.247 / 6.678 / 0.0196 / 10.582 24.77M 133.37G
+$\checkmark$ $\checkmark$ 26.443 / 6.060 / 0.0186 / 8.357 26.15M 135.83G
+Table 12: Ablation of attention in PGFAB.
+Without PSNR / SAM / RMSE / ERGAS #Params FLOPs
+Full model 26.443 / 6.060 / 0.0186 / 8.357 26.15M 135.83G
+Spat (RDG) 25.330 / 6.595 / 0.0196 / 9.894 14.37M 106.49G
+Spec (FRDB) 26.475 / 6.224 / 0.0182 / 9.812 18.96M 122.12G
+Spat &\&& Spec 25.403 / 6.575 / 0.0196 / 10.628 5.64M 96.48G
+Figure 15: PromptHSI and other All-in-One RGB image restoration methods for different degradations.
+E.2 Visualization Results Analysis
+t-SNE visulization.
+As shown in Figure 10, we compare our PromptHSI framework with other All-in-One image restoration methods to identify different degradations.
+Our approach demonstrates the superior discriminative ability for both multiple and composite degradations within t-SNE visualizations.
+This indicates that our text-prompts and feature modulation mechanism can effectively guide the model in identifying degradation types compared with conventional prompt-based All-in-One RGB image restoration methods.
+Spectral signature analysis.
+HSIs are characterized by their spectral signatures, where each substance exhibits a unique spectral profile.
+Therefore, accurate restoration in spectral domain for degraded HSIs is crucial for downstream applications, such as unmixing or material identification.
+Figure 16 presents the spectral reflectance curves between Ground-truth and reconstructed HSI, demonstrating that the proposed PromptHSI outperforms other methods in these cases.
+Figure 16: Spectral signatures across 172 bands under three different composite degradation conditions.
+Figure 17: Visual comparison of single-degradation restoration results with peer methods.
+Figure 18: Visual comparison of composite-degradation restoration results with peer methods.